# False Discovery Rate and Localizing Power

### v2

@@ 5.3 Localizing power
Even though the above issues are a direct consequence of fdr not possessing the subsetting property, i.e., not providing localizing power, routine strategies for data analysis and visualization may obscure this property. For example, the use of symmetrical thresholds is common in imaging visualization software: the user loads a statistical map and applies a threshold $t$; values above +t+t are shown in a hot color scheme (e.g., red to yellow), while values below t-t are shown in a cool color scheme (e.g., blue to cyan); for two-tailed tests, this seems appropriate and rigorous. Yet, the amount of errors in any direction is not controlled with fdr, only the global amount of errors, something that is not always apparent to the user of such tools, or they may (incorrectly) assume that the fdr is controlled in the positive and negative sides of the map. Another situation is when drawing conclusions about the amount of errors within a region after controlling globally; it might seem that controlling the error rate for the whole brain is the most rigorous, such that inferences within sub-regions would be conservative, given that a larger number of tests is taken into account. Yet, the amount of errors within any particular region is not controlled, only the global amount of errors. Any attempt to isolate a particular direction or a particular subset of tests will fail to guarantee the desired proportion of false discoveries within that subset.
@@ 5.4 More power with bky
An important yet little appreciated fact is that the bh procedure — the most commonly adopted — does not control the amount of false discoveries at the exact level $q$ desired by the researcher, but at a lower level $qV_{0}/V$ that is determined by the (unknown) proportion of tests in which the null hypothesis is true among the total number of tests (Section 2.3). Hence, the procedure is, in fact, conservative in relation to the level desired by the researcher. Despite this, it remains more powerful than procedures that control fwer such as Šidák or Bonferroni that use, for all tests, a fixed significance threshold that determines whether a p-value is significant; unless $V_{0}=V$ (complete null), $\textsc{fwer}=1$ for any fdr-controlling procedure. Whereas bh is adaptive in the sense that the threshold varies according to the sorted p-value at position $i$, with a fixed slope given by $q/V$, the bky procedure is even more adaptive than bh in that the slope itself varies according to the sorted p-value at position $i$ as q/(V+1i(1q))q/(V+1-i(1-q)) (Figure 1). The quantity q/(V+1i(1q))q/(V+1-i(1-q)) is only smaller than $q/V$ if i<1/(1q)i<1/(1-q). Thus, for typical $q$, usually well below 0.5, bky is more powerful than bh as long as smallest ($i=1$) test is rejected. As far as power considerations are concerned, the bky procedure is therefore preferable over the bh procedure.
This extra power comes with a computational price: bh runs in $\mathcal{O}(V)$ time, whereas bky runs in $\mathcal{O}(V^{2})$ time. This means that, while the time it takes to run fdr grows linearly with the number of tests to be corrected, for bky it grows quadratically. Another point worth mentioning is that bky-adjusted p-values can easily be above 1 (Figure 2). This is not a unique feature of bky: other multiple testing procedures can also lead to adjusted p-values being higher than 1, the simplest being Bonferroni, that leads to adjusted p-values that are equal to the uncorrected p-values multiplied by the number of tests $V$. This fact does not have substantial implications for hypotheses testing, and can be solved via truncation at 1.
The bky procedure was developed as an improvement over bh, and thus it would seem that the same considerations regarding dependencies between tests would hold. However, results have only been proven for the independent case (Benjamini et al., 2006), even if it tends to perform well in simulations under certain forms of dependence (e.g., Stevens et al., 2017). Scaling the thresholds of bky by $1/c(V)$ to accommodate arbitrary dependencies, as in by, while compelling, has not been proven to control the fdr either (nor that it does not).
@@ 5.5 Dependencies among tests
While we only considered one case of dependencies among tests, and observed adequate control over the error rate, existing results in the literature (Genovese et al., 2002; Logan and Rowe, 2004; Reiner-Benaim, 2007) demonstrate that, for typical cases, the modification proposed by Benjamini and Yekutieli (2001) for a more robust control over the false discovery proportion (the by procedure) is not necessary, and that even in the presence of negative correlations, bh, being more substantially more powerful, is preferable over by. That extends to bky as well.
@@ 5.6 Visualization of results
As mentioned in Section 2.6 and 5.3, localization and direction errors can be avoided by refraining from making localized or directional statements encompassing subsets of the tests, making only such statements when they refer to the whole set of tests that were subjected to the fdr procedure. Direction errors can also be avoided by using strategies such as those that we named canonical or split-tails. These approaches, however, require that the user is able to display the statistical maps in software that enables non-symmetric thresholds to be used. That is, the ideal imaging tool would accommodate separate thresholds for positive and for the negative sides of the map, such that these can be set individually after running fdr with one of these strategies. This can work also with approaches that do not use abrupt thresholds, but instead a transparency gradient, allowing subthreshold evidence be viewed informatively while highlighting suprathreshold results (Allen et al., 2012; Taylor et al., 2025).
Currently, very few software packages allow the user to easily set asymmetric thresholds for positive and negative sides of the map. Having this feature available across different imaging software would more easily allow for directional investigations with fdr and visualization of results, even using two-tailed tests. For example, a map of uncorrected p-values can be thresholded at the fdr-controlling level, or a map of adjusted p-values can be thresholded at the desired level. Alternatively, if the distribution of the test statistic is known, the threshold p-value can be converted to a threshold statistic (using the cumulative distribution function, parametric or non-parametric), which can be applied to partition statistical map into significant and non-significant regions. A thresholded map (of statistics or of p-values) can be binarized and used as a mask to display or select regions for further analyses; the same workaround can be used for multiple regions, each of them subjected to correction separately, possibly with the test level further scaled down by the number of such regions should the researcher wish to control the error rate under the complete null.
@@ 5.7 Recommendations
When applying the fdr in neuroimaging studies with two-tailed tests, researchers should adapt their approach to their goals. fdr does not ensure error control within subsets, so localized inferences require applying fdr separately, possibly with the test level adjusted to accommodate the case of no effects anywhere. For directional inferences, such as detecting positive or negative effects, the canonical or split-tails strategies are recommended to control the fdr within each direction, though the total fdr may double in the complete absence of effects; the test level is proportionally reduced if that is a concern. To apply this approach, software have to allow thresholding to be asymmetric. For non-directional inferences, the two-tailed or combined strategies suffice. The bky procedure offers greater power than the bh method.
@@ 6 Conclusion
While fdr procedures provide more power than fwer methods if many tests are expected to be significant and the user is willing to accept some false positives, caution is warranted when making directional inferences after using two-tailed tests. Strategies like running fdr separately on positive and negative test statistics, or on one-tailed p-values partitioned by hypothesis direction, avoid direction errors and control the false discovery proportion in each tail. However, these strategies lose fdr benefits under the complete null, such as controlling the fwer in these cases. Researchers should refrain from making directional inferences about fdr-corrected maps after two-tailed tests particularly when signals are not symmetrically distributed, since weak control means the proportion of false discoveries is only guaranteed globally; for the same reason, localization of effects within subsets is not possible using these methods. The bky procedure allows for additional power, albeit it does not eliminate issues with asymmetric signals. Allowing asymmetric thresholds in neuroimaging software would enable appropriate visualization after partitioned fdr.
@@
Code related to this paper is available at https://github.com/andersonwinkler/fdrfun.
@@ Acknowledgements
Support was provided by the National Institutes of Health. A.M.W.: U54-HG013247; P.A.T.: ZIC-MH002888; T.E.N.: R01-EB026859, U24-DA041123, and U19-AG073585; C.R.: RF1-MH133701 and P50-DC014664. The authors are grateful to the anonymous reviewers who provided helpful feedback.

### v3

@@ 5.4 Localizing power
Even though the issues with lack of localizing power are a direct consequence of fdr not possessing the subsetting property described in Sec. 2.6, routinely applied strategies for data analysis and visualization may obscure this fact. For example, the use of symmetrical thresholds is common in imaging visualization software: the user loads a statistical map and applies a threshold $t$; values above +t+t are shown in a hot color scheme (i.e., red to yellow), while values below t-t are shown in a cold color scheme (i.e., blue to cyan); for two-tailed tests, this seems appropriate, clear and rigorous. Yet, the amount of errors in any particular direction is not controlled with fdr, only the global amount of errors, something that is not always considered by researchers, or they may (incorrectly) assume that the fdr is controlled in the positive and negative sides of the map. Another situation is when drawing conclusions about the amount of errors within a region after controlling globally; it might seem that controlling the error rate for the whole brain is the most rigorous, such that inferences within sub-regions would be conservative, given that a larger number of tests is taken into account. Yet, the amount of errors within any particular region is not controlled, only the global amount of errors. With currently used methods, any attempt to isolate a particular direction or a particular subset of tests will fail to guarantee the desired proportion of false discoveries within that subset.
Direction and localization errors can be avoided by refraining altogether from making localized or directional statements encompassing subsets of the tests, making only such statements when they refer to the entire set subjected to fdr. Or they can be avoided rigorously with the use of selective inference, i.e., with the bb procedure. For directional inferences, those that we named canonical + bb and split-tails + bb can be used.
5.5 Real data
While simulations using synthetic data allow quantification of the observed proportion of false discoveries and how it diverges from the nominal, expected rate, those using real data allow for the visualization of the different patterns that may be observed in typical studies, depending on the fdr method and correction strategy adopted. With real data, the canonical and split-tails led to broadly similar statistically significant spatial maps; likewise, the combined and two-tailed strategies were broadly similar one to another. Two-tailed tests can be seen as a particular case of multiple testing correction, since they consider both directions of an effect. Hence, it is not surprising that the two-tailed strategy produces nearly identical results as the combined strategy, which is similar to correction across contrasts (Alberton2020).
It should be understood, however, that areas of positive signal that appeared as significant with the two-tailed and combined strategies in the narps dataset example may have emerged simply because there is a preponderance of signal in the negative side of the map; evidence for that comes from the results using synthetic data. This is because procedures that control fdr are adaptive, that is, they start with a stringent control for the smallest p-value, and proceed in steps, each contingent on the findings of the previous one, progressively loosening stringency and continuing as long as each subsequent test continue to be declared significant (Benjamini2000; Nichols2007). With two-tailed and combined strategies, the large amount of tests with strong signal in one side of the map (negative in our example) allows more opportunities for tests in the opposite side (positive) to be declared significant. The canonical and split-tails strategies eliminate the opportunity for eventual strong or widespread signals present in one side impact the other, as fdr is applied separately to each side. When used with the bb procedure, significance $t$ thresholds are further increased, mitigating the risk of excess false positives that would occur under the complete null.
@@ 5.6 More power with bky
An important yet little appreciated fact is that the bh procedure — the most commonly fdr method used — does not control the amount of false discoveries at the exact level $q$ desired by the researcher, but at a lower level $qV_{0}/V$ that is determined by the (unknown) proportion of tests in which the null hypothesis is true among the total number of tests (Section 2.3). Hence, the procedure is, in fact, conservative in relation to the level desired by the researcher. Despite this, it remains more powerful than procedures that control fwer such as Šidák or Bonferroni that use, for all tests, a fixed significance threshold that determines whether a p-value is significant; unless $V_{0}=V$ (complete null), $\textsc{fwer}=1$ for any fdr-controlling procedure. Whereas bh is adaptive in the sense that the threshold varies according to the sorted p-value at position $i$, with a fixed slope given by $q/V$, the bky procedure is even more adaptive than bh in that the slope itself varies according to the sorted p-value at position $i$ as q/(V+1i(1q))q/(V+1-i(1-q)) (Figure 1). The quantity q/(V+1i(1q))q/(V+1-i(1-q)) is only smaller than $q/V$ if i<1/(1q)i<1/(1-q). Thus, for typical $q$, usually well below 0.5, bky is more powerful than bh as long as smallest ($i=1$) test is rejected. As far as power considerations are concerned, the bky procedure is therefore preferable over the bh procedure.
This extra power comes with a computational price: bh runs in $\mathcal{O}(V)$ time, whereas bky runs in $\mathcal{O}(V^{2})$ time. This means that, while the time it takes to run fdr grows linearly with the number of tests to be corrected, for bky it grows quadratically. Another point worth mentioning is that bky-adjusted p-values can easily be above 1 (Figure 2). This is not a unique feature of bky: other multiple testing procedures can also lead to adjusted p-values being higher than 1, the simplest being Bonferroni, that leads to adjusted p-values that are equal to the uncorrected p-values multiplied by the number of tests $V$. This does not have substantial implications for hypotheses testing, and can be solved via truncating adjusted values at 1.
The bky procedure was developed as an improvement over bh, and thus it would seem that the same considerations regarding dependencies between tests would hold. However, results have only been proven for the independent case (Benjamini2006), even if it tends to perform well in simulations under certain forms of dependence (e.g., Stevens2017). Scaling the thresholds of bky by $1/c(V)$ to accommodate arbitrary dependencies, as in by, while compelling, has not been proven to control the fdr either (nor that it does not).
@@ 5.7 Dependencies among tests
While only one case of dependencies among tests was considered, which resulted in adequate control over the error rate, existing results in the literature (Genovese2002; Logan2004; Reiner-Benaim2007) demonstrate that, for typical cases, the modification proposed by Benjamini2001 for a more robust control over the false discovery proportion (the by procedure) is not necessary, and that even in the presence of negative correlations, bh, being more substantially more powerful, is preferable over by. That extends to bky as well.
The bb procedure is valid under two dependence cases: (a) when sets are independent one from another (with any dependence structure within set), provided that the procedure in the first stage is local within that set, and (b) when all p-values are prds on the subset of p-values corresponding to the true null hypotheses (Benjamini2014, Theorems 1 and 3). Nonetheless, for the combined + bb method, which clearly violates prds (since $1-p_{i}$ is fully and negatively dependent of $p_{i}$), the procedure still performed well. The reason is that, although independence among sets is violated, that violation affects only p-values that do not contribute substantially to the selection of sets: for small $p_{i}$ (approaching zero), the corresponding $1-p_{i}$ is too far from being significant to affect the outcomes from the selection rule used in the first stage; for large $p_{i}$ (around 0.5), neither $p_{i}$ or $1-p_{i}$ is close enough to typical test levels (usually around 0.05) to cause an impact, and thus the selection procedure in stage 1 remains local.
@@ 5.8 Visualization of results
The approaches that effectively control the false discovery rate in each one of the two sides of the map, or more generally in any partitioning of the map are straightforward for software developers to implement. However, their use in practice requires that the user is able to display the resulting statistical map with thresholds that are not constrained to be symmetric across both sides of the map. That is, the ideal imaging tool would accommodate separate thresholds for positive and for the negative sides of the map containg the test statistics, such that these can be set separately after running fdr. This also applies to approaches that do not use abrupt thresholds, but instead a transparency gradient, allowing subthreshold evidence be viewed informatively while highlighting suprathreshold results (Allen2012; Taylor2025).
Currently, very few software packages allow the user to easily set asymmetric thresholds for positive and negative sides of the map. Having this feature available across different imaging software would more easily allow for directional investigations with fdr and visualization of the relevant results. In the absence of software with such features, workarounds can be considered. For example, a map of uncorrected p-values can be thresholded at the fdr-controlling level, or a map of adjusted p-values can be thresholded at the desired test level, then displayed. Alternatively, if the distribution of the test statistic is known, the threshold p-value can be converted to a threshold statistic (using the cumulative distribution function, parametric or non-parametric), which can be applied to partition statistical map into significant and non-significant regions. A thresholded map (of statistics or of p-values) can be binarized and used as a mask to display or select regions for further analyses, possibly including sub-threshold results with transparency; the same workaround can be used for multiple regions, each of them subjected to correction separately, possibly with the test level further scaled down by the number of such regions should the researcher wish to control the error rate under the complete null.
@@ 5.9 Recommendations
The average error rate across sets of tests can be controlled successfully with the split-tails + bb strategy, while preventing direction errors. More generally, the main recommendation is that whenever there is interest in making statements about parts of a statistical map, the researcher should split it into disjoint sets and proceed with bb to ensure control over the average false discovery rate. A second recommendation is to use the bky procedure as routine, since it offers greater power than the bh procedure. A third recommendation concerns features in visualization software: to view results at the significance level that controls the error rate in each side (positive, negative) of the statistical map, software should ideally allow for thresholds to be asymmetric; without such feature, users need to manually mask images to view each side separately.
@@ 6 Conclusion
While fdr procedures provide more power than fwer methods if many tests are expected to be significant and the user is willing to accept some false positives, caution is needed when making directional inferences after using two-tailed tests. Using selective inference (bb), in which tests are split into sets, each subjected to local control and then all tests are corrected at an adjusted level, controls the expected average fdr, eliminates excess error rate, and generalizes to any form of localized inference. Using bky in lieu of bh as a general method to control the fdr yields greater power.
@@
Code used for assessment of the various strategies and methods is available at:
https://github.com/andersonwinkler/fdrfun.
Implementation of the methods for applied use is available in palm (Permutation Analysis of Linear Models):
https://github.com/andersonwinkler/PALM
@@ Acknowledgements
Support was provided by the National Institutes of Health. A.M.W.: R01-MH139547 and U54-HG013247; P.A.T.: ZIC-MH002888; T.E.N.: R01-EB026859, U24-DA041123, and U19-AG073585; C.R.: RF1-MH133701 and P50-DC014664. The contributions of the NIH author was made as part of his official duties as NIH federal employee, are in compliance with agency policy requirements, and are considered Works of the United States Government; however, the findings and conclusions presented in this paper are those of the authors and do not necessarily reflect the views of the NIH or the U.S. Department of Health and Human Services. The authors are grateful to the anonymous reviewers who provided helpful feedback.